OpenStudy (anonymous):

solve p^2-p+1=0 please show working

OpenStudy (anonymous):

(2p-1)(2P-1)=0 p=1/2

OpenStudy (anonymous):

how do you get p =1/2

OpenStudy (anonymous):

\[p^{2}-p+1=0\] by using the quadratic formula, \[p=\frac{ -b \pm \sqrt{b^{2} - 4ac }}{ 2a } \] where, a=coeff of p^2, b=coeff of p and c=constant term \[p=\frac{ 1 \pm \sqrt{(-1)^{2} - 4(1)(1) }}{ 2(1) }\] \[p=\frac{ 1 \pm \sqrt{1-4} }{ 2 }\] \[p=\frac{ 1 \pm \sqrt{-3} }{ 2 }\] \[\sqrt{-3}\] is an imaginary number \[\sqrt{-3} = \sqrt{3}i\] so, \[p= \frac{ 1 \pm \sqrt{3}i }{ 2 }\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

write a quadratic equation having the given solutions -2,8 please help

OpenStudy (anonymous):

here x=-2 and x=8 x=-2 x+2=0 and x=8 x-8=0 so, (x+2)(x-8)=0 x^2-6x-16=0 is the equation